OpenStudy (anonymous):

A 5 foot sign casts a shadow 8 feet long. How long a shadow does the building behind it cast? (Building is 16 feet tall)

OpenStudy (anonymous):

5/8=16/h, cross multiply 8(16)=5h 128=5h and divide by 5 answer equals 25.6